OpenStudy (anonymous):

√a -√b ------ √a + √b Rational the denominator. Assume that all expressions under radicals represent positive numbers.

OpenStudy (blacksteel):

Rationalizing a denominator means removing all roots from it. To rationalize a denominator with a sum of 2 roots, we multiply the top and bottom of the expression by the difference of the two roots: \[(\sqrt{a} - \sqrt{b}) / (\sqrt{a} + \sqrt{b}) = \] \[[\sqrt{a} - \sqrt{b} / \sqrt{a} + \sqrt{b}] *[\sqrt{a} - \sqrt{b} / \sqrt{a} - \sqrt{b}] =\] \[ (a - b - 2\sqrt{a*b} )/ (a - b)\]